The sum of two forces is 18 N and resultant whose direction is at right angles to the smaller force is 12 N.the magnitude of the forces are

Grade:12

1 Answers

Arun
6 years ago
Suppose those two forces are P & Q.
Hence according tge question
P+ Q = 18
And resultant of P &Q when they are at 90 degrees=
= 12
Now
P^2- Q^2 = 144 hence 
(P+Q) (P-Q) = 144
But P+Q = 18
Hence P-Q = 8
On solving equations 
P= 13 N
Q = 5 N